The square of the speed v of an object undergoing acceleration a is a variable of a

function a and the displacement S according to the expression given by

𝑣2 = π‘˜π‘Žπ‘šπ‘†π‘›

What dimensions should k have in order for the expression to be dimensionally

consistent?


1
Expert's answer
2021-03-09T15:28:05-0500

The units at the left hand side of the equality are:


"[v^2] = \dfrac{m^2}{s^2}"

The units at the right hand side of the equalily are:


"[ka^mS^n] = [k]\dfrac{m^m}{s^{2m}}m^n = [k]\dfrac{m^{m+n}}{s^{2m}}"

Comparing with the left hand side and expressing the dimension of "k" , obtain:


"\dfrac{m^2}{s^2} = [k]\dfrac{m^{m+n}}{s^{2m}}\\
[k] = \dfrac{m^2}{s^2}\dfrac{s^{2m}}{m^{m+n}} = m^{2-n-m}s^{2m-2}"

where "m,s" stand for meters and seconds respectively, and "m,n" in the exponents stand for numbers from the initial formula 𝑣2Β = π‘˜π‘Žπ‘šπ‘†π‘›.


Answer. "[k] = m^{2-n-m}s^{2m-2}".